OpenStudy (anonymous):

Let u = <3, -1>, v = <-6, -6>. Find 9u + 2v

OpenStudy (raffle_snaffle):

multiply <3,-1> by 9

OpenStudy (raffle_snaffle):

same goes with 2v

OpenStudy (whpalmer4):

to multiply a vector such as \(u = <3,-1>\) by a scalar, multiply the scalar times each component. For example\[3<1,2> = <3*1,3*2> = <3,6>\]

OpenStudy (anonymous):

so for <-6,-6>, it'll be <-6,-12>

zepdrix (zepdrix):

Hmm you only multiplied the second component by 2. What about the first silly? :o

OpenStudy (anonymous):

i thought you multiply it by 1 so it'll be still -6

zepdrix (zepdrix):

In palmer's example he STARTED with the 1. He multiplied each component by the `same value`, 3. In your problem, you want to multiply each component by the `same value`, 2.

OpenStudy (anonymous):

oh i see, silly me...

OpenStudy (anonymous):

so for both points it'll be <27,-9> and <-12,-12>?

zepdrix (zepdrix):

Ah very good! :) Then addition yes?

OpenStudy (anonymous):

so then 27+(-12)=15 and -9+(-12)=-21 which overall it'll be <15,-21>

zepdrix (zepdrix):

yay good job.

OpenStudy (anonymous):

thank you so much!